OpenStudy (anonymous):

Justin has 6 bags of dog food. His dogs eat 3/4 bag each day. How long will his dog food last? A. 6 3/4 days B. 8 days C. 7 days D. 4 1/2 days

OpenStudy (lukecrayonz):

B! 3/4x=6

OpenStudy (anonymous):

i think its 8 days

OpenStudy (anonymous):

6*(3/4)=4.5

OpenStudy (anonymous):

and left part is 1.5 which will be completed in 2 more days...i.e. 3/4 +3/4

OpenStudy (anonymous):

i twill be 8days

Directrix (directrix):

@madrockz Just wondering - why didn't you divided 6 by (3/4) and get 8 directly? 6/(3/4) = 6 * (4/3) = 2*4 = 8?.

OpenStudy (anonymous):

it was my way of solving! ;)

Directrix (directrix):

Got it, mental arithmetic. Cool.

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

B